Ten firms compete in a market to sell product X. The total sales of all firms selling the
product are $2 million. Ranking the firms’ sales from highest to lowest, we find the
top four firms’ sales to be $260,000, $220,000, $150,000, and $130,000, respectively.
Calculate the four-firm concentration ratio in the market for product X. (LO1)
Answer #1

Total sales by four firms = 260000 + 220000 + 150000 + 130000 = 760000

Four firm concentration ratio = Total sales by four firms / total sales by all firms = 760000 / 2000000 = 0.38 or 38%